# Reexamining Technological Support for Genealogy Research, Collaboration, and Education

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer-Supported Cooperative Work and Social Computing; 2024; Costa Ricaccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 2.1. The Genealogy Research Process
Academic studies of genealogy research practices and community have been scattered across various disciplines, including information science (Yakel, 2004; Fulton, 2005; Mansourian, 2021), library and archival science (Duff and Johnson, 2003; Carini, 2016), sociology (Lambert, 2002; Kramer, 2011; Bottero, 2015), tourism (Kennedy-Eden and Gretzel, 2021; Prince, 2022; Urrestarazu-Capellán et al., 2022), and public history (Groot, 2020; Evans, 2020). In 2003, Duff and Johnson (2003) presented a pioneering user study focusing on genealogists’ information seeking behavior with archival systems. Their work revealed the complexity of genealogical research, which is more than just “collecting names and building family trees”. They found that genealogists need to refer to historical context and apply searching strategies in their research. As archival systems did not meet their needs, genealogists developed their own finding aids, consulted others who had knowledge about a period, geographic area, or community, and formed a strong network through “courses, genealogical society meetings, and newsletters” (Duff and Johnson, 2003).
@@
$2.$ Sensemaking Challenges and Information Literacy
@@
As genealogy research comprises extensive use of archives and libraries, another research field closely related is information literacy, a theoretical framework describing an individual’s capability to “seek, evaluate, use and create information (United Nations Educational, Scientific and Cultural Organization, 2023).” This concept was originated from the library and information science discipline. Though traditionally focused on print culture, it evolved in past years — associated with the concept of digital and media literacy, which emphasize aspects working with digital information and computer — to address information challenges today (of College and Libraries, 2000; Mackey and Jacobson, 2011; Belshaw et al., 2012). Information literacy skills were recognized as a critical component in combating online misinformation (Khan and Idris, 2019; Sharon and Baram-Tsabari, 2020; Jones-Jang et al., 2021). In the genealogy domain, librarians and archivists focused on educating genealogy beginners frequently encounter audiences with limited information and digital literacy skills (Bremer, 2018; Wilkinson and Bond, 2021). They reflected on amateurs’ insufficiency with computer systems and accessing online information, underlining the importance of investigating how amateurs’ skills could be improved in a technology-intensive genealogy research context.
@@ 3.2. Interview Design
$3.$ Data Collection and Analysis
@@ 4.2.2. Keeping track with research logs
A lot of genealogists, professionally, use some kind of a log to track their searches. …And I just keep track of what I looked at, what source I consulted, and then I write down what was my purpose and intention for the search, what was I looking for. And then I also track what were the specific criteria that I used. So if it’s a database, “Did I search for all people with just the first name?” “Did I look for all people with the last name?” “Did I use a wildcard to give me spelling variation?” So I’$l$ keep track of the search parameters, just so that I know how thorough my search was, and then I’$l$ make notes about the outcome. “Did I find some potential candidates?” “Did I find the ancestor I wanted?” Or maybe I didn’t. So I keep track of all of that, and I make comments. — P9E
@@
I’m the person that’$l$ have 30 tabs open, and I’$l$ be opening tabs until it forces me to stop, or, you know, I’ve got a little notebook next to me that I’m making weird gibberish notes. And then if I go back a week later, I’m like “I don’t know what the hell this means!” — P15A
@@ 4.3.2. Verifying clues by oneself
Even though expert genealogists are willing to leverage research hints and others’ trees in their research, they are also selective when evaluating if one piece of information can serve as evidence in formulating their conclusion. In general, experts expressed their preference of using primary information from original records as evidence, while considering the rest, such as others’ family trees and algorithmic research hints, as research clues that they need to verify themselves. In P6E’s words, “I’$l$ take it [someone’s tree] as a hint. And I’$l$ do what I have to do to validate it.”
@@
There’s no way most people have enough records, enough proof of who 6,000 people are to put them in there. …it’$l$ show, say, 6,000 people, 6,500 records. So you go, “You have barely over one record per person. ’Cause I have people with 20’s and 30’s [records per person], that tells me that you have not clearly looked at their stuff.” — P19A
@@
Participants further acknowledged that such behavior is common among amateurs. As P18A elaborated, “There’s a big problem in genealogy community at most of the beginner’s stage …where I’$l$ just copy information over and over.” Considering how easy for users to accept research hints and duplicate others’ family trees on genealogy websites, one can quickly expand his/her family tree by replicating unverified information. However, expert genealogists emphasized that the frequency of one piece of information does not imply its accuracy. P1E argued, ”Just because it’s there 40 times doesn’t mean it’s true.”
$4.$ GPS Point 4: Resolution of Conflicting Evidence
@@ 4.6.1. Privacy concerns in online collaboration
I’$l$ send my students there [a Facebook group] because it’s really really helpful. If I have a question, people can answer it, and there’s experts there that will answer and they do it very quickly. So it’s a great way to get help, and with little cost or no cost. — P1E
We heard about diverse practices for sharing genealogy information from our participants. For some amateurs, their enthusiasm in sharing genealogical information goes beyond genealogy-focused websites, extending to general purpose websites like social media. Several participants expressed concern for the risk of revealing personal information in genealogy research, who may limit their appearance on social media and provide just enough information to an online community when asking for help. It may make the online collaboration less effective, but their privacy concerns are not baseless. Government records are usually released to the public decades after they are generated for privacy reasons (Heimlich, 2012). Genealogy websites follows a similar rule, limiting access to profiles on living people in the tree (Ancestry, [n. d.], 2023). In the meantime, more than one amateur participant mentioned researching living individuals who they think might be their relatives using public online information:
@@ 5.3. Computer Supported Andragogy
Individuals’ genealogy research often starts as a hobby, and we found amateurs’ learning experience fits in well with the concept of andragogy, i.e., adult education and learning, as opposed to pedagogy, the education of dependent personalities such as children and students (Knowles, [n. d.]; Forrest III and Peterson, 2006). Andragogy is characterized with problem-oriented, just-in-time learning, and leverages experience and knowledge of learners (Forrest III and Peterson, 2006). Indeed, we found that many amateurs prefer learning during research. Their learning is self-directed, motivated by the challenge they are facing, and they need to identify the right learning resource that is suitable for their particular knowledge level.

### v3

@@ Abstract.
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 2article: CSCW155publicationmonth: 4price: doi: 10.1145/3711053ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 2.1. The Genealogy Research Process
Academic studies of genealogy research practices and community have been scattered across various disciplines, including information science (Yakel, 2004; Fulton, 2005; Mansourian, 2021), library and archival science (Duff and Johnson, 2003; Carini, 2016), sociology (Lambert, 2002; Kramer, 2011; Bottero, 2015), tourism (Kennedy-Eden and Gretzel, 2021; Prince, 2022; Urrestarazu-Capellán et al., 2022), and public history (Groot, 2020; Evans, 2020). In 2003, Duff and Johnson (2003) presented a pioneering user study focusing on genealogists’ information seeking behavior with archival systems. Their work revealed the complexity of genealogical research, which is more than just “collecting names and building family trees.” They found that genealogists need to refer to historical context and apply searching strategies in their research. As archival systems did not meet their needs, genealogists developed their own finding aids, consulted others who had knowledge about a period, geographic area, or community, and formed a strong network through “courses, genealogical society meetings, and newsletters” (Duff and Johnson, 2003).
@@
2.2. Sensemaking Challenges and Information Literacy
@@
As genealogy research comprises extensive use of archives and libraries, another research field closely related is information literacy, a theoretical framework describing an individual’s capability to “seek, evaluate, use and create information” (United Nations Educational, Scientific and Cultural Organization, 2023). This concept was originated from the library and information science discipline. Though traditionally focused on print culture, it evolved in past years — associated with the concept of digital and media literacy, which emphasize aspects working with digital information and computer — to address information challenges today (of College and Libraries, 2000; Mackey and Jacobson, 2011; Belshaw et al., 2012). Information literacy skills were recognized as a critical component in combating online misinformation (Khan and Idris, 2019; Sharon and Baram-Tsabari, 2020; Jones-Jang et al., 2021). In the genealogy domain, librarians and archivists focused on educating genealogy beginners frequently encounter audiences with limited information and digital literacy skills (Bremer, 2018; Wilkinson and Bond, 2021). They reflected on amateurs’ insufficiency with computer systems and accessing online information, underlining the importance of investigating how amateurs’ skills could be improved in a technology-intensive genealogy research context.
@@ 3.2. Interview Design
3.3. Data Collection and Analysis
@@ 4.2.2. Keeping track with research logs
A lot of genealogists, professionally, use some kind of a log to track their searches. …And I just keep track of what I looked at, what source I consulted, and then I write down what was my purpose and intention for the search, what was I looking for. And then I also track what were the specific criteria that I used. So if it’s a database, “Did I search for all people with just the first name?” “Did I look for all people with the last name?” “Did I use a wildcard to give me spelling variation?” So I’ll keep track of the search parameters, just so that I know how thorough my search was, and then I’ll make notes about the outcome. “Did I find some potential candidates?” “Did I find the ancestor I wanted?” Or maybe I didn’t. So I keep track of all of that, and I make comments. — P9E
@@
I’m the person that’ll have 30 tabs open, and I’ll be opening tabs until it forces me to stop, or, you know, I’ve got a little notebook next to me that I’m making weird gibberish notes. And then if I go back a week later, I’m like “I don’t know what the hell this means!” — P15A
@@ 4.3.2. Verifying clues by oneself
Even though expert genealogists are willing to leverage research hints and others’ trees in their research, they are also selective when evaluating if one piece of information can serve as evidence in formulating their conclusion. In general, experts expressed their preference of using primary information from original records as evidence, while considering the rest, such as others’ family trees and algorithmic research hints, as research clues that they need to verify themselves. In P6E’s words, “I’ll take it [someone’s tree] as a hint. And I’ll do what I have to do to validate it.”
@@
There’s no way most people have enough records, enough proof of who 6,000 people are to put them in there. …it’ll show, say, 6,000 people, 6,500 records. So you go, “You have barely over one record per person. ’Cause I have people with 20’s and 30’s [records per person], that tells me that you have not clearly looked at their stuff.” — P19A
@@
Participants further acknowledged that such behavior is common among amateurs. As P18A elaborated, “There’s a big problem in genealogy community at most of the beginner’s stage …where I’ll just copy information over and over.” Considering how easy for users to accept research hints and duplicate others’ family trees on genealogy websites, one can quickly expand his/her family tree by replicating unverified information. However, expert genealogists emphasized that the frequency of one piece of information does not imply its accuracy. P1E argued, ”Just because it’s there 40 times doesn’t mean it’s true.”
4.4. GPS Point 4: Resolution of Conflicting Evidence
@@ 4.6.1. Privacy concerns in online collaboration
I’ll send my students there [a Facebook group] because it’s really really helpful. If I have a question, people can answer it, and there’s experts there that will answer and they do it very quickly. So it’s a great way to get help, and with little cost or no cost. — P1E
We heard about diverse practices for sharing genealogy information from our participants. For some amateurs, their enthusiasm in sharing genealogical information goes beyond genealogy-focused websites, extending to general purpose websites like social media. Several participants expressed concern for the risk of revealing personal information in genealogy research, who may limit their appearance on social media and provide just enough information to an online community when asking for help. It may make the online collaboration less effective, but their privacy concerns are not baseless. Government records are usually released to the public decades after they are generated for privacy reasons (Heimlich, 2012). Genealogy websites follows a similar rule, limiting access to profiles on living people in the tree (Ancestry, 2023b, a). In the meantime, more than one amateur participant mentioned researching living individuals who they think might be their relatives using public online information:
@@ 5.3. Computer Supported Andragogy
Individuals’ genealogy research often starts as a hobby, and we found amateurs’ learning experience fits in well with the concept of andragogy, i.e., adult education and learning, as opposed to pedagogy, the education of dependent personalities such as children and students (Knowles, 1973; Forrest III and Peterson, 2006). Andragogy is characterized with problem-oriented, just-in-time learning, and leverages experience and knowledge of learners (Forrest III and Peterson, 2006). Indeed, we found that many amateurs prefer learning during research. Their learning is self-directed, motivated by the challenge they are facing, and they need to identify the right learning resource that is suitable for their particular knowledge level.